OpenStudy (anonymous):

Point N on the number line shows Rob's score after the first round of a quiz: In round two, he lost 6 points. Which expression shows how many total points he has at the end of round two? 1 + (−6) = −5, because −5 is 6 units to the left of 1 1 + 6 = −5, because −5 is 6 units to the left of 1 1 + (−5) = −6, because −6 is 5 units to the left of 1 1 + 5 = −6, because −6 is 5 units to the left

OpenStudy (anonymous):

This is the number line

OpenStudy (texaschic101):

1 + (-6) = 1 - 6 = -5 ....because -5 is 6 units to the left of 1......it is the first one